OpenStudy (anonymous):

A vehicle uses 1 tablespoon of gasoline to drive 125 yards. How many miles can the vehicle travel per gallon?

OpenStudy (tkhunny):

Multiply by 1. \(\dfrac{1\;tablespoon}{125\;yds}\cdot\dfrac{1\;cup}{16\;tablespoon}\cdot\dfrac{1\;yd}{3\;ft}\cdot...\) Find more factors that equal one (1) and multiply by them until you get the units you want. 2 cups = 1 pint 2 pints = 1 quart 4 quarts = 1 gallon 5280 ft = 1 mile You'll need them all unless youhappen to remember some other conversions that will get you there.

OpenStudy (anonymous):

hi..can you help me with this question: a vehicle uses 1 tablespoon of gasoline to drive 125 yards. how many miles can the vehicle travel per gallon?

OpenStudy (wolf1728):

1 gallon = 256 tablespoons So for 1 gallon you can travel 256*125 yards = 32,000 yards= 96,000 feet = 18.1818 ... miles